OpenStudy (anonymous):

Which points are 24 units from the point located at –51 on a number line? A. –75 and 75 B. –75 and –27 C. –27 and 27 D. –27 and 75

OpenStudy (anonymous):

You add 24 to -51 to get the point on the right You subtract 24 from -51 to get the point on the left.

OpenStudy (anonymous):

so whats the answer

OpenStudy (anonymous):

-51 + 24 = -27 -51 - 24 = -75 What do you think it's the answr along with this equation

OpenStudy (anonymous):

D

OpenStudy (anonymous):

nope....look at it closely

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

-27 and -75 is....

OpenStudy (anonymous):

btw it's not a

OpenStudy (anonymous):

B

OpenStudy (anonymous):

CORRECT!